You planted a 16-inch hemlock tree that grows at a rate of 4 inches per year. Then 5 years later you planted a 10-inch blue spruce tree that grows at a rate of 6 inches per year. In how many years after you planted the hemlock tree will the two trees be the same heigh? How tall will the trees be?
Answer by scott8148(5895)

let x=years

"You planted a 16-inch hemlock tree that grows at a rate of 4 inches per year"
__ 16+4x

"5 years later you planted a 10-inch blue spruce tree that grows at a rate of 6 inches per year"
__ 10+6(x-5)

"how many years after you planted the hemlock tree will the two trees be the same heigh?
__ 16+4x=10+6(x-5)

16+4x=10+6x-30 __ 36=2x

substitute value of x into 16+4x to find height